# Measuring software innovation with open source software development data

### v1

@@ 3.2.3 Data Analysis
To explore software releases’ effect on future utilization, we calculated log-differences in dependents from release date to one year after release. We select a one-year period because it is typically sufficient for users to integrate new features, adapt to changes, and evaluate implications of a release on their project [58, 59, 60, 61]. These log-differences are a measure of how much the software’s dependency network expanded during this period. We categorized these measurements by release type—major, minor, and patch—and conducted an Analysis of Variance (ANOVA) to determine whether there were statistically significant differences between groups. We also performed post-hoc pairwise Welch’s independent t-tests to identify which specific release types exhibited significantly different distributions in the one-year change in dependents. Both the ANOVA and t-tests were executed using the SciPy library [68].
Next, we compared log-differences in dependents for prevalent ecosystems in our dataset: JavaScript, Python, and Ruby. This comparison allowed us to observe variations in software utilization across these dominant programming ecosystems for understanding ecosystem-specific behaviors. Further stratification of the dataset was conducted to investigate how dependent growth changes when controlling for version series (i.e., zero-version, one-version, and two-plus-version) and initial software utilization. For this, we divided the initial number of dependents into three groups: bottom-25th percentile, middle-50th percentile, and top-25th percentile. By employing these subgroup analyses, we aimed to reveal how both software maturity and early-stage adoption influence long-term dependency growth. Data visualizations were generated using seaborn [69], matplotlib [70], opinionated [71], and colormaps [72].
@@ 3.3.1 Data Generation
For the analysis of GitHub releases and complexity, we closely follow the methodology established in § 3.2. We combine multiple private daily snapshot datasets encompassing repository information, package releases, and dependency details, ensuring consistency through uniform filtering and processing techniques. We perform two more steps to enrich our dataset. First, we focus on package releases that include an accompanying GitHub release, which allows us to capture the associated release notes. Second, we implement a stringent filtering criterion, removing releases lacking a minimum of 512 text characters in their release notes. This requirement ensures that the releases included in our dataset contain sufficient textual information, enabling both our prompted LLM and domain experts to effectively assess the complexity of each release. These notes potentially provide critical context and insights into the changes made in each release.
@@ 3.3.2 Data Processing
Our data processing methodology involved a novel and structured approach to assess the complexity of GitHub releases. After compiling a dataset of release notes and repository metadata, we moved to the task of rating the complexity of each release to better understand the effort required to implement the listed changes.
@@
Given that this complexity-scoring approach is novel, we thoroughly examined its robustness. To validate the model’s output, we compared its ratings with those provided by professional developers. A JavaScript developer evaluated the complexity of releases from JavaScript projects, while a Python developer did the same for Python projects. This human annotation process served as a benchmark to assess the accuracy of the GPT-4 model’s ratings. Our analysis revealed a positive correlation between the model and the developers’ assessments, with a correlation coefficient of $\alpha$ = 0.39 for JavaScript (p ¡ 0.00001) and $\alpha$ = 0.36 for Python (p ¡ 0.001). Notably, more than 60% of the GPT-4 and developer ratings were within one complexity rank of each other (JavaScript: 68.5%, Python: 63.3%), indicating a strong degree of alignment between the model and human experts.
@@
Despite this upward bias, the model demonstrated a strong degree of agreement with the professional developers’ ratings within one complexity rank, particularly at the median and higher complexity levels. Given this high level of concordance, we proceeded to use the finalized GPT-4 prompt to generate complexity ratings for the entire dataset of GitHub Releases and their corresponding release notes. This approach allowed us to efficiently scale the complexity assessment process while maintaining reasonable accuracy and alignment with expert human judgment.
@@ 3.3.3 Data Analysis
Python 12157 4.52 1.16 4 5 5
JavaScript 11288 3.90 1.28 3 4 5
Ruby 2605 4.27 1.08 4 4 5
To evaluate differences in the distributions of complexity ratings between ecosystems, we conducted an ANOVA test. We then performed post-hoc pairwise Welch’s independent t-tests to identify where significant differences occurred between specific ecosystems. Finally, we sought to understand how complexity ratings were related to the subsequent software utilization by calculating the Spearman correlation ($\rho$) between the complexity rating of each release and the one-year log-difference in dependents. This analysis was performed separately for each major ecosystem in our dataset. The goal was to assess whether more complex releases tend to result in greater or lesser growth in the number of dependents over time, providing insight into how release complexity may influence the broader software ecosystem’s uptake and reliance on the software in question. ionResults In this section, we present the results of our analysis. We begin by reporting the results about dependency measures in § 3.4. Subsequently, we report the results on release complexity as annotated by LLMs in § 3.5.
@@ 3.4 Repository metrics and package dependencies
3.4.1 Major releases associated with significant increases in package dependencies
Key finding: At the one-year post-release mark, major releases are associated with significantly larger increases in package dependencies compared to minor or patch releases across the JavaScript, Python, and Ruby ecosystems, highlighting distinct patterns of dependency growth that emerge following significant package updates.
We find that packages experience a statistically significant positive change in their log-difference in direct dependents following a Major release compared to other release types (see Figure 2). Table 4 provides a detailed overview of the log-difference in direct dependents associated with different release types across various programming languages. Specifically, we observe that JavaScript major releases are associated with an average increase of 2.32 in log-difference in dependents, while minor and patch releases show comparatively modest increases of only +0.15 and +0.16, respectively. In the Python ecosystem, major releases correspond to an average increase of 1.79 in log-difference in dependents, in stark contrast to the minimal increases of +0.11 for minor releases and +0.08 for patch releases. Similarly, Ruby major releases demonstrate a strong impact, linked to an average increase of 2.43 in log-difference in dependents, whereas minor and patch releases exhibit negligible increases of +0.02 and +0.08, respectively.
Overall, in this one-year interval, these results indicate that major releases significantly enhance the log-difference in direct dependents across all three ecosystems, with JavaScript exhibiting the most pronounced effect. This underscores the critical role of major updates in driving package adoption and shaping dependency growth patterns in the short-term.
Figure 2: Log-Difference in Dependents after One-Year Post Software Release Split by Software Release Type
Table 4: Log-Difference in Dependents after One-Year Post Software Released Split by Language
Average Increase in Log-Difference in Dependents JavaScript Python Ruby
Major release 2.32 1.79 2.43
Minor release 0.15 0.11 0.02
Patch release 0.16 0.08 0.08
Figure 3: Distributions of Log-Difference in Dependents after 180, 360, 540, and 720 days since Software Release. Each row of plots are all the same ecosystem (JavaScript, Python, Ruby), while each column of plots are all the same initial package utilization bins (bottom 25th percentile, middle 50th percentile, and top 25th percentile).
3.4.2 Consistent dependency growth patterns across ecosystems and time frames
Key finding: Major releases generate larger positive changes in log-difference compared to minor and patch releases, regardless of the package ecosystem, initial package utilization, or time frame examined up to two years post-release. However, the magnitude of the log-difference tends to diminish slightly in the second year.
Next, we examine log-difference dynamics in direct package dependencies over a two year time frame. We observe that the trend of major releases maintaining a strong positive impact on dependency growth persists across all ecosystems during this period (see Figure 3). However, it is important to note that the magnitude of the log-difference tends to diminish slightly in the second year, in particular for the Ruby ecosystem, where we observe a strong decline in log-difference for major releases by the third year, resulting in significantly lower values compared to the other ecosystems. This suggests a potential saturation effect or a shift in package utilization patterns specific to Ruby. The Python ecosystem also experiences a reduction in log-difference over time, indicating a gradual decrease in the impact of major releases. Meanwhile, the JavaScript ecosystem demonstrates remarkable stability, with log-difference values remaining relatively consistent, highlighting its resilience in maintaining dependency growth from major releases.
@@ 3.5.1 Correlation between release complexity and dependency growth
Key finding: Release complexity, as rated by a prompted GPT-4 model, correlates slightly positively with dependency growth, particularly for JavaScript and Python packages. However, the Ruby ecosystem does not exhibit the same correlation, suggesting different dynamics in the Ruby community and a different handling of release notes.
We observe statistically significant differences in complexity ratings across the three main release types: major, minor, and patch. Major releases consistently show the highest complexity, as they typically introduce significant changes or breaking features. Patch releases, on the other hand, exhibit the lowest complexity ratings, reflecting their focus on bug fixes or minor adjustments. Minor releases lie between these two. This pattern holds consistent across all three ecosystems—JavaScript, Python, and Ruby. As depicted in Figure 4(a), major releases display the highest complexity ratings, followed by minor and then patch releases, which is according to our expectations and serves as a robustness check of our analysis. The differences in complexity between these release types are statistically significant and can be attributed to the inherent nature of the changes introduced in each type. See Table 5 for a summary of the statistical analysis, including means, standard deviations, and p-values for significance testing across ecosystems.
@@
JavaScript major minor 4.83 (±1.07) 4.13 (±1.03) ¡ 0.0001
patch 4.83 (±1.07) 3.38 (±1.18) ¡ 0.0001
minor patch 4.13 (±1.03) 3.38 (±1.18) ¡ 0.0001
Python major minor 5.19 (±1.23) 4.76 (±1.0) ¡ 0.01
patch 5.19 (±1.23) 4.16 (±1.14) ¡ 0.0001
minor patch 4.76 (±1.0) 4.16 (±1.14) ¡ 0.0001
Ruby major minor 5.52 (±1.02) 4.54 (±0.87) ¡ 0.0001
patch 5.52 (±1.02) 3.93 (±1.03) ¡ 0.0001
minor patch 4.54 (±0.87) 3.93 (±1.03) ¡ 0.0001
Next, we explore the relationship between release complexity and package dependency growth. The JavaScript and Python ecosystems hold statistically significant, slightly positive correlations between the complexity of releases and the log-difference in dependents after one year (see Figure 4(b)). Specifically, for JavaScript package releases, we observe a Spearman’s correlation coefficient of 0.13 ($p\leq 0.00001$), suggesting that more complex releases tend to have marginally higher growth in dependent packages over time. Similarly, Python package releases show a Spearman’s correlation coefficient of 0.08 ($p\leq 0.0001$), reflecting a similar trend, albeit weaker than in JavaScript. Curiously, the Ruby ecosystem does not exhibit the same statistical significance between release complexity and dependency growth. This points to different dynamics in the Ruby community and a different handling of release notes, leading to a divergence in how complexity is documented and perceived. It is also possible that other factors, such as community practices or the characteristics of dependencies, play a more significant role in influencing package adoption in Ruby.

### v2

@@ 3.2.3 Data Analysis
To explore the software releases’ effect on future utilization, we calculated log-differences in dependents from the release date to one year after release. We select a one-year period because it is typically sufficient for users to integrate new features, adapt to changes, and evaluate implications of a release on their project [58, 59, 60, 61]. These log-differences are a measure of how much the software’s overall dependency network expanded during this period. We categorized these measurements by release type—major, minor, and patch—and conducted an Analysis of Variance (ANOVA) to determine whether there were statistically significant differences between groups. We also performed post-hoc pairwise Welch’s independent t-tests to identify which specific release types exhibited significantly different distributions in the one-year change in dependents. The ANOVA and t-tests were executed using the SciPy library [68].
Next, we compared log-differences in dependents for prevalent ecosystems in our dataset: JavaScript, Python, and Ruby. This comparison allowed us to observe variations in software utilization across these dominant programming ecosystems for understanding ecosystem-specific behaviors. Further stratification of the dataset was conducted to investigate how dependent growth changes when controlling for version series (i.e., zero-version, one-version, and two-plus-version) and initial software utilization. For this, we divided the initial number of dependents into four groups: less than 100 dependents (small), at least 100 but less than 1,000 dependents (medium), at least 1,000 but less than 10,000 dependents (large), and at least 10,000 dependents (huge). We employed these subgroup analyses to reveal how software maturity and early-stage adoption influence long-term dependency growth. Data visualizations were generated using seaborn [69], matplotlib [70], opinionated [71], and colormaps [72].
@@ 3.3.1 Data Generation
We closely follow the methodology established in § 3.2 for the analysis of GitHub releases and complexity. We combine multiple private daily snapshot datasets encompassing repository information, package releases, and dependency details, ensuring consistency through uniform filtering and processing techniques. We perform two more steps to enrich our dataset. First, we focus on package releases that include an accompanying GitHub release, which allows us to capture the associated release notes. Second, we implement a stringent filtering criterion, removing releases lacking a minimum of 512 text characters in their release notes. This requirement ensures that the releases in our dataset contain sufficient textual information, enabling both our prompted LLM and domain experts to effectively assess the complexity of each release. These notes potentially provide critical context and insights into the changes made in each release.
@@ 3.3.2 Data Processing
Our data processing methodology involved a novel and structured approach to assessing the complexity of GitHub releases. After compiling a dataset of release notes and repository metadata, we rated the complexity of each release to better understand the effort required to implement the listed changes.
@@
Given that this complexity-scoring approach is novel, we thoroughly examined its robustness. To validate the model’s output, we compared its ratings with those provided by professional developers. A JavaScript developer evaluated the complexity of releases from JavaScript projects, while a Python developer did the same for Python projects. This human annotation process served as a benchmark to assess the accuracy of the GPT-4 model’s ratings. Our analysis revealed a positive correlation between the model and the developers’ assessments, with a correlation coefficient of $\alpha$ = 0.39 for JavaScript (p $<$0.00001) and $\alpha$ = 0.36 for Python (p $<$0.001). Notably, more than 60% of the GPT-4 and developer ratings were within one complexity rank of each other (JavaScript: 68.5%, Python: 63.3%), indicating substantial alignment between the model and human experts.
@@
Despite this upward bias, the model demonstrated substantial agreement with the professional developers’ ratings within one complexity rank, particularly at the median and higher complexity levels. Given this high level of concordance, we used the finalized GPT-4 prompt to generate complexity ratings for the entire dataset of GitHub Releases and their corresponding release notes. This approach allowed us to efficiently scale the complexity assessment process while maintaining reasonable accuracy and alignment with expert human judgment.
@@ 3.3.3 Data Analysis
Python 12186 4.52 1.16 4 5 5
JavaScript 11469 3.90 1.28 3 4 5
Ruby 2606 4.27 1.08 4 4 5
We conducted an ANOVA test to evaluate differences in the distributions of complexity ratings between ecosystems. We then performed post-hoc pairwise Welch’s independent t-tests to identify where significant differences occurred between specific ecosystems. Finally, we sought to understand how complexity ratings were related to the subsequent software utilization by calculating the Spearman correlation ($\rho$) between the complexity rating of each release and the one-year log-difference in dependents. This analysis was performed separately for each major ecosystem in our dataset. The goal was to assess whether more complex releases tend to result in greater or lesser growth in the number of dependents over time, providing insight into how release complexity may influence the broader software ecosystem’s uptake and reliance on the software in question.
4 Results
In this section, we present the results of our analysis. We begin by reporting the results about dependency measures in § 4.1. Subsequently, we report the results on release complexity as annotated by LLMs in § 4.2.
@@ 4.1 Repository metrics and package dependencies
4.1.1 Release adoption patterns are ecosystem and pre-existing utilization specific
Key finding: Semantic versioning types exhibit ecosystem-specific and pre-release utilization-dependent patterns in predicting dependency growth at both one-year and two-year post-release timeframes. Minor releases demonstrate relatively consistent adoption benefits across JavaScript, Python, and Ruby ecosystems and initial dependency levels. In contrast, major and patch releases vary significantly based on ecosystem culture and baseline package popularity.
Across the studied ecosystems and package sizes, the impact of release type on the log-difference in direct dependents one year after release reveals nuanced patterns (Table 4 and Figure 2, see Appendix B and C for multiple-timepoint distributions over one and two years post-release). In the NPM ecosystem, we observe that major and minor releases tend to correlate with the largest average increase in log-difference for small, medium, and large-sized packages, respectively.
The PyPI ecosystem exhibits a different dynamic. Small PyPI packages show the largest average growth in direct dependents following patch releases, while medium and large-sized packages see comparable increases across all release types. Interestingly, major and minor releases are associated with the most substantial log-difference increase for huge PyPI packages.
In contrast, the RubyGems ecosystem shows a generally muted response to new releases regarding dependent growth. While small and medium-sized packages show slightly higher average log-differences after major and minor releases, respectively, the overall increases are not statistically significantly different. Notably, the largest RubyGems packages show virtually no positive change or slight negative trends for major and patch releases. This may reflect broader trends in the adoption and utilization of the Ruby ecosystem during the study period.
Table 4: Mean and Standard Deviation of Log-Difference in Direct Dependents One Year After Release, Stratified by Ecosystem, Package Size, and Release Type. Values in bold indicate the statistically significantly highest mean log-difference within each ecosystem and package size category when comparing major, minor, and patch releases (pairwise t-tests).
Ecosystem Package Size Log Diff. Major Log Diff. Minor Log Diff. Patch
npm small 0.104 $\pm$ 0.422 0.12 $\pm$ 0.368 0.107 $\pm$ 0.384
medium 0.117 $\pm$ 0.336 0.103 $\pm$ 0.253 0.103 $\pm$ 0.246
large 0.104 $\pm$ 0.209 0.081 $\pm$ 0.176 0.09 $\pm$ 0.18
huge 0.042 $\pm$ 0.073 0.044 $\pm$ 0.098 0.044 $\pm$ 0.094
pypi small 0.07 $\pm$ 0.226 0.088 $\pm$ 0.242 0.114 $\pm$ 0.273
medium 0.088 $\pm$ 0.144 0.092 $\pm$ 0.135 0.088 $\pm$ 0.143
large 0.063 $\pm$ 0.094 0.076 $\pm$ 0.107 0.069 $\pm$ 0.101
huge 0.043 $\pm$ 0.04 0.044 $\pm$ 0.038 0.033 $\pm$ 0.017
rubygems small 0.087 $\pm$ 0.399 0.081 $\pm$ 0.333 0.063 $\pm$ 0.315
medium 0.05 $\pm$ 0.168 0.065 $\pm$ 0.186 0.059 $\pm$ 0.204
large 0.012 $\pm$ 0.063 0.023 $\pm$ 0.067 0.015 $\pm$ 0.046
huge -0.002 $\pm$ 0.017 0.0 $\pm$ 0.017 -0.001 $\pm$ 0.019
Figure 2: Mean Log-Difference in Direct Dependents One Year After Release Across Ecosystems and Package Sizes. This set of heatmaps illustrates the average log-difference in direct dependents for packages in the NPM, PyPI, and RubyGems ecosystems, categorized by package size (small, medium, large, huge) and release type (Major, Minor, Patch). The color intensity, consistent across all subplots, represents the magnitude of the mean log-difference, allowing for direct visual comparison of growth patterns across different ecosystems and package sizes following a new version release.
These findings suggest that the impact of a new package version on its adoption, as measured by the growth in direct dependents, is not uniform across ecosystems or package sizes. This variability highlights the importance of considering these factors when analyzing package impact, innovation, and ecosystem health.
4.1.2 Release adoption patterns are ecosystem and version-series specific
Key finding: Analysis across versioning series maturity (zero-ver, one-ver, two-plus-ver packages) reveals that ecosystem-specific adoption patterns persist regardless of development lifecycle stage, with minor releases consistently performing as a stable middle ground across all versioning series, reinforcing their role as reliable innovation signals.
Building on our understanding of how pre-existing package utilization influences adoption, we now examine the interplay between semantic versioning types and a package’s version series maturity (zero-ver, one-ver, and two-plus-ver) in predicting dependent growth. This analysis, detailed in Table 5 and Figure 3 (See Appendix D and E for multiple-timepoint distributions over one and two years post-release, indicates that ecosystem-specific adoption patterns largely persist regardless of a package’s development lifecycle stage. In the NPM ecosystem, major releases correlate with increased mean log-difference in direct dependents for both zero-ver (0.219 $\pm$ 0.525) and one-ver (0.281 $\pm$ 0.724) packages. For two-plus-ver NPM packages, mean log-difference values across major, minor, and patch releases are comparable.
The PyPI ecosystem presents a different dynamic. Major releases show a higher mean log-difference for zero-ver packages (0.101 $\pm$ 0.247). One-ver PyPI packages show comparable growth across release types. For two-plus-ver PyPI packages, patch releases show a higher mean log-difference (0.123 $\pm$ 0.265). The RubyGems ecosystem, consistent with previous findings, generally exhibits lower dependent growth across all version series. Minor releases show a higher mean log-difference for zero-ver RubyGems packages (0.15 $\pm$ 0.435), while major releases are higher for one-ver packages (0.136 $\pm$ 0.526). Differences in dependent growth across release types are minimal and low for two-plus-ver RubyGems packages.
Table 5: Mean and Standard Deviation of Log-Difference in Direct Dependents One Year After Release, Stratified by Ecosystem, Version Series, and Release Type. Values in bold indicate the statistically significantly highest mean log-difference within each ecosystem and package size category when comparing major, minor, and patch releases (pairwise t-tests).
Ecosystem Version Series Log Diff. Major Log Diff. Minor Log Diff. Patch
npm zero-ver 0.219 $\pm$ 0.525 0.204 $\pm$ 0.549
one-ver 0.281 $\pm$ 0.724 0.152 $\pm$ 0.396 0.136 $\pm$ 0.416
two-plus-ver 0.086 $\pm$ 0.32 0.092 $\pm$ 0.285 0.089 $\pm$ 0.304
pypi zero-ver 0.101 $\pm$ 0.247 0.086 $\pm$ 0.259
one-ver 0.099 $\pm$ 0.266 0.091 $\pm$ 0.237 0.08 $\pm$ 0.224
two-plus-ver 0.065 $\pm$ 0.191 0.086 $\pm$ 0.209 0.123 $\pm$ 0.265
rubygems zero-ver 0.125 $\pm$ 0.399 0.15 $\pm$ 0.435
one-ver 0.136 $\pm$ 0.526 0.07 $\pm$ 0.288 0.061 $\pm$ 0.317
two-plus-ver 0.05 $\pm$ 0.257 0.06 $\pm$ 0.254 0.047 $\pm$ 0.22
Figure 3: Mean Log-Difference in Direct Dependents One Year After Release Across Ecosystems and Package Sizes. This set of heatmaps illustrates the average log-difference in direct dependents for packages in the NPM, PyPI, and RubyGems ecosystems, categorized by version series (zero-version, one-version, two-plus-version) and release type (Major, Minor, Patch). The color intensity, consistent across all subplots, represents the magnitude of the mean log-difference, allowing for direct visual comparison of growth patterns across different ecosystems and version series following a new version release.
These results indicate that semantic versioning’s influence on package adoption remains connected to ecosystem culture, even as packages mature.
@@ 4.2.1 Correlation between release complexity and dependency growth
Key finding: A release’s technical complexity correlates with semantic versioning type but shows minimal direct correlation with dependency growth, indicating that versioning serves as a signaling mechanism that drives adoption independently of actual technical complexity
Next, we investigated the ”technical complexity to implement the changes” within a release, as rated by GPT-4, to understand how release complexity relates to semantic versioning types and subsequent package adoption. Our analysis, presented in Table 6 (and Figure 4(a)), indicates a consistent pattern across JavaScript, Python, and Ruby: Major releases exhibit higher mean complexity ratings than Minor releases, which have higher mean ratings than Patch releases. For instance, in JavaScript, Major releases had a mean complexity of 4.71 ($\pm$1.10), Minor releases 4.14 ($\pm$1.00), and Patch releases 3.39 ($\pm$1.17). All pairwise comparisons between each language’s release types show statistically significant differences (p $<$0.0001). This correlation between semantic versioning type and assessed release complexity suggests that maintainers generally adhere to semantic versioning conventions regarding the scope and effort involved in their releases.
@@
JavaScript major minor 4.71 ($\pm$1.10) 4.14 ($\pm$1.00) $<$0.0001
patch 4.71 ($\pm$1.10) 3.39 ($\pm$1.17) $<$0.0001
minor patch 4.14 ($\pm$1.00) 3.39 ($\pm$1.17) $<$0.0001
Python major minor 5.34 ($\pm$1.06) 4.76 ($\pm$1.02) $<$0.0001
patch 5.34 ($\pm$1.06) 4.17 ($\pm$1.15) $<$0.0001
minor patch 4.76 ($\pm$1.02) 4.17 ($\pm$1.15) $<$0.0001
Ruby major minor 5.09 ($\pm$1.05) 4.55 ($\pm$0.87) $<$0.0001
patch 5.09 ($\pm$1.05) 3.96 ($\pm$1.03) $<$0.0001
minor patch 4.55 ($\pm$0.87) 3.96 ($\pm$1.03) $<$0.0001
However, a different trend is observed when examining the relationship between this assessed technical complexity and downstream adoption, measured by the log difference in dependents one year post-release. As shown in Figure 4(b), each programming language’s distributions of log differences across complexity ratings (1 to 7) indicate limited correlation. Spearman’s rank correlation coefficients for log-difference in dependents versus complexity rating are low for JavaScript (statistic = 0.015, p = 0.220) and Python (statistic = 0.024, p = 0.118), indicating no significant relationship in these ecosystems. For Ruby, a weak negative correlation is present (statistic = -0.057, p = 0.024), suggesting that higher complexity ratings are associated with slightly lower dependent growth. This general absence of a strong or consistent positive correlation indicates that the internal technical challenges or scale of changes involved in a release, from a maintainer’s perspective, do not appear to be a key factor influencing whether end-users upgrade or integrate a new package version.
@@
The combined insights from these analyses offer a dual perspective. Maintainers appear to align their internal development efforts with semantic versioning conventions, producing releases whose technical complexity generally corresponds to their declared version type (Major, Minor, or Patch). However, the limited correlation between release complexity and dependent growth indicates that downstream users’ decisions to adopt new package versions are likely driven by factors other than the maintainers’ development effort. This implies that end-users prioritize external factors such as new functionality, bug fixes, or perceived value rather than the intrinsic difficulty of the release when considering an upgrade.